# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ III-E Positional Encoding
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}+L\left(\tilde{\textit{{V}}}\right)\right]\approx\\ &\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right.\right.\\ &\left.\left.+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$ (17)
where $L(\cdot)$ denotes a depthwise convolution operation with a kernel size of 3. Fig. 1 shows the comparison of the vanilla attention mechanism and ELFATT. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$.
III-F Complexity Analysis
According to Eqs. (16) and (17), the complexity of ELFATT can be divided into two parts: (a) the global linear attention head ${\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right)$; (b) the local blockify attention head $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)$. The complexity of the global linear attention head is $O(m\times c_{1}^{2})$ in the case of $m>c_{1}$, and $O(m^{2}\times c_{1})$ in the case of $m\leq c_{1}$. The complexity of the local blockify attention head is $O((m^{2}/b)\times c_{2})$. If $m/b\ll m$, the local blockify attention head will also achieve almost linear complexity.
IV Experiments and Results
We evaluated ELFATT in commonly used vision tasks: image classification (ImageNet-1K [42]), semantic segmentation (ADE20K [43]), and object detection (MS COCO 2017 [44]). We compared ELFATT with the vanilla softmax-based attention mechanism [38], the memory-efficient attention mechanism (FlashAttention-2 [7]), local window-based attention mechanisms (Swin [22] and CSWin [23]), and kernel-based attention mechanisms (Agent [19], EFFATT [18], and FLatten [10]). The backbone ViT architectures used to evaluate the different attention mechanisms are: Swin-T [22], Swin-B [22], CSWin-T24181 [10], and CSWin-B36292 [10]. The original Swin-T, CSWin-T24181, and CSWin-B36292 are called Swin-T-LOCAL, Swin-B-LOCAL, CSWin-T24181-LOCAL, and CSWin-B36292-LOCAL, respectively. The backbones after replacing local window-based attention mechanisms with the vanilla softmax-based attention mechanism [38] are called Swin-T-GLOBAL, Swin-B-GLOBAL, CSWin-T24181-GLOBAL, and CSWin-B36292-GLOBAL, respectively. The backbones after replacing local window-based attention mechanisms with kernel-based attention mechanisms (Agent [19], EFFATT [18], and FLatten [10]) are called Swin-T-Agent, Swin-T-EFFATT, Swin-T-FLatten, Swin-B-Agent, Swin-B-EFFATT, Swin-B-FLatten, CSWin-T24181-Agent, CSWin-T24181-EFFATT, CSWin-T24181-FLatten, CSWin-B36292-Agent, CSWin-B36292-EFFATT, and CSWin-B36292-FLatten, respectively. The backbones after replacing local window-based attention mechanisms with ELFATT are called Swin-T-ELFATT, Swin-B-ELFATT, CSWin-T24181-ELFATT, and CSWin-B36292-ELFATT, respectively. For the ImageNet-1K image classification task, we used the same training settings and data augmentation methods from [23] to train all models from scratch using mixed precision. For the ADE20K semantic segmentation task [43] and the MS COCO 2017 object detection task [44], we used the same training settings and data augmentation methods from [45] to fine-tune the weights of all models obtained from the ImageNet-1K image classification task using mixed precision. We also compared ELFATT with ConvNeXt-T [46] and VMamba-T [45] on ImageNet-1K, ADE20K, and MS COCO 2017. The experiments of ImageNet-1K, ADE20K, and MS COCO 2017 were carried out on 8 NVIDIA vGPU (32 GB) GPUs. Inference throughput comparison experiments were carried out on 1 NVIDIA H20 (96 GB) GPU. Besides high-performance computing applications, we also compared ELFATT with state-of-the-art attention mechanisms on edge GPUs (NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano GPU). We used FlashAttention-2 [7] to speed up all models that are compatible with FlashAttention. The PyTorch implementation of ELFATT is available at [https://github.com/Alicewithrabbit/ELFATT].
IV-A Image Classification Performance
Table I shows the performance comparison of different attention mechanisms on ImageNet-1K. From Table I, the vanilla softmax-based attention mechanism (Swin-T-GLOBAL and CSWin-T24181-GLOBAL) can outperform most linear attention mechanisms using the same architecture. Only Agent and ELFATT achieve the most close performance compared to the vanilla softmax-based attention mechanism in this paper. As to the inference speed comparison, the proposed ELFATT achieves the highest inference throughput (frame per second, FPS) than all other attention mechanisms when using CSWin-T24181 as the backbone. ELFATT achieves almost the same speed as EFFATT and is significantly faster than other attention mechanisms when using Swin-T as the backbone. ELFATT offers almost a 2x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2. The vanilla softmax-based attention mechanism using FlashAttention-2 is still 0.1-0.2x slower than ELFATT without using FlashAttention-2. With the use of FlashAttention-2 to optimize memory operations, ELFATT can be further accelerated and offers 1.2-1.3x speedups over the vanilla softmax-based attention mechanism. Figs. 2-3 show the visual comparison of different attention mechanisms using CSWin-T24181 as the backbone. ELFATT shows a much closer attention map to the vanilla softmax-based attention mechanism than other attention mechanisms. For the larger backbone, CSWin-B36292, ELFATT still achieves state-of-the-art (SOTA) performance. ELFATT offers a 2.1/1.4x speedup over the vanilla softmax-based attention mechanism with/without using FlashAttention-2 for an input resolution of $224^{2}$ and a 3.3/1.8x speedup over the vanilla softmax-based attention mechanism with/without using FlashAttention-2 for an input resolution of $384^{2}$. In addition, ELFATT achieves significant leading performance under the backbone of Swin-B, offering a 1.8x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 for an input resolution of $224^{2}$ and a 2.8x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 for an input resolution of $384^{2}$. After using FlashAttention-2, it matches the speed of EFFATT which is a real linear attention mechanism. Also, as shown in Table I, it can be seen that although some methods have lower FLOPs than ELFATT, their actual speed is slower than ELFATT which is consistent with the observations and conclusions of [47].
TABLE I: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of different attention mechanisms on ImageNet-1K. Note: Inference throughput of tiny models is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU and inference throughput of base models is obtained using a batch size of 256 for $224^{2}$ resolution and 32 for $384^{2}$ resolution with mixed precision on a single NVIDIA H20 (96 GB) GPU. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-B36292-Agent $224^{2}$ 84.7 930/994 imgs/s 73M 14.92G/14.49G
CSWin-B36292-EFFATT $224^{2}$ 84.4 985/1059 imgs/s 73M 14.98G/14.53G
CSWin-B36292-ELFATT $224^{2}$ 84.7 1000/1187 imgs/s 73M 15.47G/14.46G
CSWin-B36292-FLatten $224^{2}$ 84.5 814/864 imgs/s 75M 14.96G/14.52G
CSWin-B36292-GLOBAL $224^{2}$ 84.7 478/879 imgs/s 73M 22.33G/14.39G
CSWin-B36292-LOCAL $224^{2}$ 84.4 941/1037 imgs/s 73M 15.03G/14.39G
CSWin-B36292-Agent $384^{2}$ 85.8 246/276 imgs/s 73M 46.33G/42.57G
CSWin-B36292-EFFATT $384^{2}$ 85.7 294/331 imgs/s 73M 46.57G/42.69G
CSWin-B36292-ELFATT $384^{2}$ 85.8 272/355 imgs/s 73M 51.21G/42.48G
CSWin-B36292-FLatten $384^{2}$ 85.5 238/266 imgs/s 78M 46.43G/42.67G
CSWin-B36292-GLOBAL $384^{2}$ 85.9 82/201 imgs/s 73M 110.89G/42.28G
CSWin-B36292-LOCAL $384^{2}$ 85.5 276/323 imgs/s 73M 47.06G/42.28G
CSWin-T24181-Agent $224^{2}$ 83.1 2297/2425 imgs/s 20M 4.31G/4.14G
CSWin-T24181-EFFATT $224^{2}$ 82.6 2394/2526 imgs/s 20M 4.35G/4.17G
CSWin-T24181-ELFATT $224^{2}$ 83.1 2603/2856 imgs/s 20M 4.44G/4.13G
CSWin-T24181-FLatten $224^{2}$ 83.1 1934/2025 imgs/s 21M 4.34G/4.16G
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
CSWin-T24181-LOCAL $224^{2}$ 82.7 2330/2519 imgs/s 20M 4.36G/4.09G
Swin-B-Agent $224^{2}$ 84.0 1367/ —   imgs/s 88M 15.44G/   —
Swin-B-EFFATT $224^{2}$ 84.2 1439/1536 imgs/s 88M 15.69G/15.33G
Swin-B-ELFATT $224^{2}$ 84.5 1314/1497 imgs/s 91M 16.46G/15.68G
Swin-B-FLatten $224^{2}$ 83.8 1226/ —   imgs/s 89M 15.46G/   —
Swin-B-GLOBAL $224^{2}$ 84.2 743/1325 imgs/s 88M 21.57G/15.19G
Swin-B-LOCAL $224^{2}$ 83.5 1351/ —   imgs/s 88M 15.47G/   —
Swin-B-Agent $384^{2}$ 84.9 372/ — imgs/s 88M 46.34G/   —
Swin-B-EFFATT $384^{2}$ 85.3 433/481 img/s 88M 48.18G/45.04G
Swin-B-ELFATT $384^{2}$ 85.5 372/457 img/s 91M 52.79G/46.08G
Swin-B-FLatten $384^{2}$ 85.0 353/ — imgs/s 91M 46.49G/   —
Swin-B-GLOBAL $384^{2}$ 85.3 134/313 img/s 88M 99.76G/44.64G
Swin-B-LOCAL $384^{2}$ 84.5 357/ — imgs/s 88M 47.19G/   —
Swin-T-Agent $224^{2}$ 82.6 2847/ —   imgs/s 29M 4.53G/ —
Swin-T-EFFATT $224^{2}$ 82.1 3165/3282 imgs/s 28M 4.55G/4.45G
Swin-T-ELFATT $224^{2}$ 82.7 2884/3159 imgs/s 30M 4.99G/4.67G
Swin-T-FLatten $224^{2}$ 82.1 2502/ —   imgs/s 29M 4.50G/ —
Swin-T-GLOBAL $224^{2}$ 82.4 1269/2571 imgs/s 28M 8.81G/4.38G
Swin-T-LOCAL $224^{2}$ 81.4 2943/ —   imgs/s 28M 4.51G/ —
ConvNeXt-T $224^{2}$ 82.1 3911/ —   imgs/s 29M 4.47G/ —
VMamba-T $224^{2}$ 82.5 1837/ —   imgs/s 30M 4.84G/ —
@@ IV-A Image Classification Performance
TABLE II: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different block sizes in each level of CSWin-T24181 on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
@@
IV-B Ablation Study of Block Size
To validate the effect of the block size on the performance of ELFATT, we performed an ablation study using different block sizes at each level of CSWin-T24181-ELFATT on ImageNet-1K. As shown in Table II, with the increasing number of block sizes, the inference speed of ELFATT will slow down, while the classification accuracy of ELFATT shows an increasing trend. ELFATT using 49-49-196-49 and 196-196-196-49 blocks achieve the best performance in terms of speed and accuracy. It seems that FlashAttention-2 acceleration is more efficient for some block sizes (with FlashAttention-2 acceleration, the block size of 196 is faster than the block size of 49 which may be caused by GPU architectures). Hence, the block sizes for ELFATT can be determined according to the performance and efficiency requirements.
IV-C Ablation Study of LePE
To validate the effect of LePE on performance, we compared the performance of ELFATT and the vanilla softmax-based attention mechanism using or without using LePE in CSWin-T24181 on ImageNet-1K. As shown in Table III, using LePE, the performance of the vanilla softmax-based attention mechanism improves from 82.9 to 83.1 and the performance of ELFATT also improves from 82.9 to 83.1, respectively, further demonstrating the powerful local positional information enhancement of LePE. More details about LePE can be found in [23].
TABLE III: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT and the vanilla softmax-based attention mechanism using or without using LePE in CSWin-T24181 on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “w/o” denotes “without”, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
@@ IV-C Ablation Study of LePE
CSWin-T24181-ELFATT $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
CSWin-T24181-ELFATT-w/o-LePE $224^{2}$ 82.9 2801/3271 imgs/s 20M 4.54G/4.12G
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
CSWin-T24181-GLOBAL-w/o-LePE $224^{2}$ 82.9 1332/2296 imgs/s 20M 7.58G/4.08G
TABLE IV: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by using the ELFATT modules to replace the vanilla softmax-based attention modules in different levels of CSWin-T24181-GLOBAL on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration. ✓denotes the level consists of full ELFATT modules and ✔ denotes half of this level is composed of ELFATT modules and the other half is composed of vanilla softmax-based attention (VaniATT) modules. The order of composition is: VaniATT-ELFATT-...-VaniATT-ELFATT.
Levels with using the ELFATT module Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
Level 1 Level 2 Level 3 Level 4
@@
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
TABLE V: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by CSWin-T24181-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
@@
IV-D Ablation Study of Number of Levels Using the ELFATT Module

### v4

@@ Appendix A Approximation Error Bound Analysis
$\begin{split}&\left\lVert\left[\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top},\ \mathrm{exp}(\tilde{\textit{{Q}}})\mathrm{exp}(\tilde{\textit{{K}}})^{\top}\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}})\mathrm{exp}(\tilde{\textit{{K}}})^{\top}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
and for LOCAL is as
$\begin{split}&\left\lVert\left[\left(\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\odot\textit{{Z}}_{1}\right),\ \left(\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}_{2}\right)\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\odot\textit{{Z}}_{1}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}_{2}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
where $\xi=F$ denotes the Frobenius norm, and $\textit{{Z}}_{1}$ and $\textit{{Z}}_{2}$ are different to Z of ELFATT, because LOCAL uses a more complex cross-shaped blockify method for different heads. All upper bounds were obtained from the final attention layer of the second level of the backbone. The upper bounds of different methods show a decreasing trend as the training process progresses. ELFATT further reduces the upper bound of the LOCAL attention mechanism, although it still has a higher bound than EFFATT, which is caused by its sparse blockify attention heads. Figs. 2 (b) and (c) show the comparison of relative attention matrix approximation error ($\left\lVert\textit{{A}}-\textit{{A}}^{{}^{\prime}}\right\rVert_{\xi}/\left\lVert\textit{{A}}\right\rVert_{\xi}$, where A denotes the attention matrix of VaniATT and $\textit{{A}}^{\prime}$ denotes the attention matrix of linear attention) of global linear attention heads of ELFATT and the corresponding heads in EFFATT and relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the corresponding heads in LOCAL. The backbone used is CSWin-T (Han et al., 2023). Since ELFATT uses a hybrid head architecture, its half heads consist of global linear attention heads, and the remaining heads consist of sparse blockify attention heads. We compared the relative attention matrix approximation error of global linear attention heads of ELFATT and the heads in the same position of EFFATT. We compared the relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the heads in the same position of LOCAL. We found that the global linear attention heads of ELFATT have a lower approximation error than the heads in the same position of EFFATT, and the sparse blockify attention heads of ELFATT have a lower approximation error than the heads in the same position of LOCAL. The hybrid head architecture has complementary effects on approximation error reduction.
(a) Upper Bounds
(b) Global Linear Attention Heads
(c) Sparse Blockify Attention Heads
Figure 2. (a) The upper bound of different attention mechanisms (EFFATT (Shen et al., 2021), ELFATT, and LOCAL (Dong et al., 2022)) for approximating attention matrices of vanilla attention during the training of ImageNet-1K. (b) The comparison of relative attention matrix approximation error of global linear attention heads of ELFATT and the corresponding heads in EFFATT. (c) The comparison of relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the corresponding heads in LOCAL. The backbone used is CSWin-T.
(a) True Label: American Staffordshire Terrier
(b) True Label: Hartebeest
(c) True Label: Zebra
(d) True Label: Library
(e) True Label: Snail
(f) True Label: Black-Footed Ferret
(g) True Label: Scottish Deerhound
(h) True Label: Pineapple
Figure 3. Visual comparison of class activation map (CAM) based attention results of different attention mechanisms obtained using Score-CAM (Wang et al., 2020b). Note: Backbone used is CSWin-T (Han et al., 2023).
(a) CSWin
(b) Swin
Figure 4. Accuracy-efficiency curves obtained by different attention mechanisms on ImageNet-1K using different bacbones (CSWin (Han et al., 2023) and Swin (Liu et al., 2021)). Efficiency is represented by the inference throughput of each method. All methods’ inference throughput and accuracy results are from Table 1 of the main body of the paper.
Appendix B Ablation Study of Different Block Sizes
To validate the effect of the block size on the performance of ELFATT, we performed an ablation study using different block sizes at each level of CSWin-T-ELFATT on ImageNet-1K. As shown in Table 9, with the increasing number of block sizes, the inference speed of ELFATT will slow down, while the classification accuracy of ELFATT shows an increasing trend. ELFATT using 49-49-196-49 and 196-196-196-49 blocks achieve the best performance in terms of speed and accuracy. It seems that FlashAttention-2 is more efficient for some block sizes (with FlashAttention-2, the block size of 196 is faster than the block size of 49 which may be caused by GPU architectures). Hence, the block sizes for ELFATT can be determined according to the performance and efficiency requirements.
Appendix C Ablation Study of LePE
To validate the effect of LePE on performance, we compared the performance of ELFATT and VaniATT using or without using LePE in CSWin-T on ImageNet-1K. As shown in Table 10, using LePE, the performance of VaniATT improves from 82.9 to 83.1 and the performance of ELFATT also improves from 82.9 to 83.1, respectively, further demonstrating the powerful local positional information enhancement of LePE. More details about LePE can be found in (Dong et al., 2022).
@@ Appendix C Ablation Study of LePE
Table 9. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different block sizes in each level (L1-L4) of CSWin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold.
@@
Table 10. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT and VaniATT using or without using LePE in CSWin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “w/o” denotes “without”, “nFA” denotes non-FlashAttention-2, and “FA” denotes FlashAttention-2.
@@
CSWin-T-ELFATT
LePE $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
w/o-LePE $224^{2}$ 82.9 2801/3271 imgs/s 20M 4.54G/4.12G
CSWin-T-GLOBAL
LePE $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
w/o-LePE $224^{2}$ 82.9 1332/2296 imgs/s 20M 7.58G/4.08G
Table 11. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by using the ELFATT modules to replace the VaniATT modules in different levels (L1-L4) of CSWin-T-GLOBAL on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2, and “FA” denotes FlashAttention-2. ✓denotes the level consists of full ELFATT modules and ✔ denotes half of this level is composed of ELFATT modules and the remaining is composed of VaniATT modules. The order of composition is: VaniATT-ELFATT-...-VaniATT-ELFATT.
Levels using ELFATT Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
L1 L2 L3 L4
@@
CSWin-T-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
Table 12. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by CSWin-T-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold.
@@
Appendix D Ablation Study of Number of Levels Using the ELFATT Module